Atlantis Fisheries issues zero coupon bonds on the market at a price of $419 per bond. If these bonds are callable in 5 years at a call price of $547, what is their yield to call? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Yield to call ____
Answer #1

This question requires application of basic time value of money function, according to which FV = PV * (1 + r) where FV is F

$547 = $419 * (1 + r)5

1.305489 = (1 + r)5

1 + r = 1.0548

YTC, r = 5.48%